# Cross-Modal Few-Shot Learning: a Generative Transfer Learning Framework

### v1

@@ Abstract
Most existing studies on few-shot learning focus on unimodal settings, where models are trained to generalize on unseen data using only a small number of labeled examples from the same modality. However, real-world data are inherently multi-modal, and unimodal approaches limit the practical applications of few-shot learning. To address this gap, this paper introduces the Cross-modal Few-Shot Learning (CFSL) task, which aims to recognize instances from multiple modalities when only a few labeled examples are available. This task presents additional challenges compared to classical few-shot learning due to the distinct visual characteristics and structural properties unique to each modality.
To tackle these challenges, we propose a Generative Transfer Learning (GTL) framework consisting of two stages: the first stage involves training on abundant unimodal data, and the second stage focuses on transfer learning to adapt to novel data. Our GTL framework jointly estimates the latent shared concept across modalities and in-modality disturbance in both stages, while freezing the generative module during the transfer phase to maintain the stability of the learned representations and prevent overfitting to the limited multi-modal samples. Our finds demonstrate that GTL has superior performance compared to state-of-the-art methods across four distinct multi-modal datasets: Sketchy, TU-Berlin, Mask1K, and SKSF-A. Additionally, the results suggest that the model can estimate latent concepts from vast unimodal data and generalize these concepts to unseen modalities using only a limited number of available samples, much like human cognitive processes.
@@ 1 Introduction
Collecting large amounts of labeled data in real-world applications is often prohibitively expensive, time-consuming, or simply impractical (Tharwat & Schenck, 2023; Sheng et al., 2024). Few-shot learning (FSL) has emerged as a viable solution, enabling models to generalize effectively using only a handful of labeled examples (Song et al., 2023; Chen et al., 2019; Luo et al., 2023; Ke et al., 2024). However, existing few-shot methods face significant challenges in handling the increasing prevalence of multi-modal data, such as multi-spectral images or multimedia content, which differs significantly from the extensive RGB data typically used in research. Multi-modal data, collected from multiple different types of sources and modalities such as different sensors or imaging protocols, is becoming increasingly essential in applications like surveillance (Wu et al., 2024; Hu et al., 2022) and medical image analysis (Jiang et al., 2023b; Mok et al., 2024). These challenges highlight the need for more advanced FSL frameworks capable of leveraging the complementary information inherent in multi-modal data (Luo et al., 2023; Jiang et al., 2023a).
Recent efforts have leveraged large pre-trained foundation models to extend their capabilities to novel multi-modal tasks, including tabular data (Ye et al., 2024), audio (Lin et al., 2023b; Duan et al., 2024), and video classification (Qing et al., 2023), moving beyond traditional unimodal tasks like image classification (Conti et al., 2023). Despite these advancements, the transfer of visual knowledge across different visual modalities remains relatively underexplored. Visual data, such as images and videos, constitute the most commonly studied data types; however, other visual modalities, such as infrared, depth, and sketches, exhibit both shared characteristics and distinct differences. These modalities share structural and contextual similarities with RGB data but possess unique attributes that make data collection and model adaptation more challenging.
To address these challenges, this paper introduces a new Cross-modal Few-Shot Learning (CFSL) task, which aims to classify instances with multiple visual modalities using only a few labeled samples per class. In this task, the multi-modal data are organized into a support set and a query set. The support set contains a small number of labeled multi-modal examples per class, serving as the basis for learning. The query set includes unlabeled instances from the same classes and modalities that need to be classified. The ultimate goal is to train a model solely on the support set that can accurately classify the query set, regardless of modality. As illustrated in Figures 1(a-c), unlike classical supervised learning and FSL confined to a single modality, CFSL incorporates data from multiple visual modalities. The primary challenge of CFSL arises from the inherent variability and domain gaps between different visual modalities, which complicates feature extraction and alignment. Additionally, the task requires models to identify and leverage shared underlying semantics across modalities while adapting to the unique characteristics of each.
@@
Figure 1: Comparison of recognition tasks. (a) Classical recognition requires extensive labeled data within a single modality. (b) Few-shot recognition uses a few labeled samples in a single modality to classify unseen samples. (c) Our proposed CFSL involves few labeled multi-modal samples and aims to generalize well to recognizing unseen multi-modal samples from the same classes, leveraging both seen and unseen data from different modalities. (d) Visual concept learning illustrates the ability to generalize concepts like “Bull” across real images and sketches (Pablo Picasso. The Bull, 1945.).
@@
Previous studies have shown that compact visual representations are closely linked to real-world images, emphasizing the importance of underlying concepts in enabling humans to identify target objects (Vinker et al., 2022; 2023; Mukherjee et al., 2024). As illustrated in Figure 1(d), these fundamental concepts can be readily learned with only a few examples (Lake & Piantadosi, 2020), thanks to the humans’ ability to generalize from limited data, regardless of visual modality. This capability, referred to as “evocation”, is grounded in the vast reservoir of previously encountered visual experiences that support recognition and generalization. Inspired by this observation, we hypothesize that models can achieve similar evocative capabilities by learning latent concepts from abundant unimodal data.
@@
We propose a Generative Transfer Learning (GTL) framework to facilitate knowledge transfer between unimodal and multi-modal data. In the GTL framework, we posit that the latent concepts underlying target objects consist of two components: (1) an intrinsic concept that captures the core characteristic shared across modalities, and (2) an in-modality disturbance that accounts for variations unique to each modality. Our approach aims to estimate these latent components and encode the relationships between the intrinsic concepts and visual content. By disentangling the intrinsic concepts from the in-modality disturbances, the GTL framework enables adapting to multi-modal data while preserving the transferable relationships learned from unimodal data, thereby improving both adaptability and accuracy across various modalities.
@@
Within the GTL framework, our methodology consists of a two-stage training process. In the first stage, the generative learning stage, the model learns latent concepts label-free, relying solely on pre-trained visual representations. This stage focuses on capturing the intrinsic concept and the variations in visual content across modalities, ensuring that the learned relationship between the latent concept and visual content is robust and transferable. In the second stage, the recognition stage, the backbone network is frozen, and a separate classifier is trained on top of the learned latent intrinsic concepts to perform label prediction. This two-stage framework enables the model to disentangle intrinsic visual concepts from in-modality disturbances, facilitating effective recognition across visual modalities.
@@
Our contributions are summarised as follows: (i) We introduce a new cross-modal few-shot learning task, which focuses on the connection and distinction between visual modalities, requiring models to perform recognition on multi-modal data with minimal labeled samples. This task better reflects real-world scenarios, where multi-modal data is scarce and diverse. (ii) We propose the generative transfer learning framework, designed to disentangle intrinsic concepts from in-modality disturbances, enabling efficient knowledge adaption across modalities. (iii) We demonstrate the effectiveness of our method through extensive experiments on multiple cross-modal datasets, including  Sketchy (Sangkloy et al., 2016)TU-Berlin (Eitz et al., 2012), the fine-grained biometric dataset  Mask1K (Lin et al., 2023a) and SKSF-A (Yun et al., 2024).
@@ 2 Task Settings
  • A support set $D_{\text{support}}=\{(x^{i}_{m},y^{i})|m\in\{m^{1},m^{2},\dots,m^{d}\},y^{i}\in Y_{\text{novel}}\}$, containing a limited number of labeled samples per class from multiple modalities. Typically, this consists of $K$ labeled samples for each of $N$ classes.
  • A query set $D_{\text{query}}=\{x^{j}_{m}|m\in\{m^{1},m^{2},\dots,m^{d}\}\}$, containing a larger number of unlabeled samples from the same classes as $D_{\text{support}}$, but potentially from the different modalities.
The fundamental challenge in this task is the limited labeled data in $D_{\text{support}}$, combined with the multi-modal nature of the data in $D_{\text{novel}}$. The scarcity of labeled data per class, along with the need to generalize across different modalities, makes this task significantly more challenging.
@@
(a) t-SNE of feature distribution of different modalities.
(b) The data generating model for CFSL.
Figure 2: The observation of the severe modality differences and the details of the proposed generative model. (a) Illustration of the modality difference by the t-SNE clustering of the pre-trained CLIP (Radford et al., 2021) features of different modalities from Mask1K dataset (Lin et al., 2023a). (b) The proposed generative process for the representation learning stage, the green symbols are assumed to be parameters that enable the models to adapt from base to novel data.
@@ 3.1 Problem Formulation
$p(\mathbf{x},\mathbf{z},\mathbf{u},y)=p_{\delta}(\mathbf{z}_{m}|\mathbf{u}^{1},\mathbf{u}^{2},\dots,\mathbf{u}^{d})p_{\gamma}(\mathbf{z}_{c}|y)p_{\theta}(\mathbf{x}|\mathbf{z}_{m},\mathbf{z}_{c})p(\mathbf{u})p(y).$ (1)
@@
$\mathbf{z}_{m}\sim p_{\delta}(\mathbf{z}_{m}|\mathbf{u}^{1},\mathbf{u}^{2},\dots,\mathbf{u}^{d})$ (4)
@@
$\mathbf{z}_{c}\sim p_{\gamma}(\mathbf{z}_{c}|y)$ (5)
@@
$\mathbf{x}\sim p_{\theta}(\mathbf{x}|\mathbf{z}_{m},\mathbf{z}_{c})$ (6)
@@ 3.2 Network Design
In this section, we introduce the key components of our network, each playing a distinct role in modeling the cross-modal few-shot classification task.
Figure 3: The proposed GTL framework. During the training on base data, all modules are trained (as in the blue dashed box), but when adapting to novel data, the generator is frozen, and all other parts are tunable (as in the red dashed box). The classifier for recognition is separately initialized on base and novel training since there are non-overlap classes between them.
@@
$q_{\alpha}(\mathbf{{z}}_{c},\mathbf{{z}}_{m}|\mathbf{x})\sim\mathcal{N}(\mu,\sigma^{2})$ (8)
@@
The in-modality disturbance latent variable $\mathbf{z_{m}}$ is assumed to be transferable when adapting from the base to novel data. Since there is a lack of direct supervision regarding which representation corresponds to modality-specific information, we employ a flexible estimator to approximate the unobservable prior $p(\mathbf{z}_{m})$. Inspired by the latent domain learning method (Deecke et al., 2022), we use a set of learnable gating functions $g(\mathbf{x})$ that assign each observation $\mathbf{x}$ to multiple latent domains, capturing different perspectives.
@@
$\mathbf{\hat{z}}_{m}^{{}^{\prime}}=h_{\delta,\mathbf{\hat{u}}\sim q_{\eta}}(\mathbf{\hat{u}},\mathbf{\hat{z}}_{m})\quad\text{and}\quad q_{\eta}(\mathbf{\hat{u}}|\mathbf{x})=\sum_{d=1}^{D}g_{d}(\mathbf{x})V_{d}(\mathbf{x}),$ (9)
@@
The reconstruction module is responsible for generating an estimate of the observation $\mathbf{\hat{x}}$ based on the estimated latent variables $\mathbf{\hat{z}_{c}}$ (class-relevant information) and $\mathbf{\hat{z}_{m}^{{}^{\prime}}}$ (modality-relevant information). We adopt a generator with a structure similar to, but reverse of, the posterior estimator. While the posterior estimator encodes the latent variables, the generator decodes them back into the observed data.
@@
$\hat{\mathbf{x}}=\text{Dropout}\left(\text{ReLU}\left(\text{MLP}(\mathbf{\hat{z}{{}^{\prime}}})\right)\right)\quad\text{where}\quad\mathbf{\hat{z}}{{}^{\prime}}=\{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}\}$ (10)
@@
The classifier $p_{\phi}(\mathbf{y}|\mathbf{z}_{c})$, parameterized by $\phi$, can be implemented using a simple linear classifier. This involves a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times c}$, where $d$ is the dimension of the latent variable $\mathbf{\hat{z}}_{c}$, and $c$ denotes the number of output classes. The classifier is trained using a standard cross-entropy loss function. The linear classifier computes the logits directly by linearly combining the latent variable $\mathbf{\hat{z}}_{c}$ with the weight matrix $\mathbf{W}$, followed by a softmax function to output class probabilities. The logits are computed as:
@@
For detailed network architecture, please refer to Appendix C.
@@ 3.3 Learning Objectives
$\mathcal{L}_{\text{ELBO}}=\underbrace{\mathbb{E}_{q_{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}|\mathbf{x}}}\left[\ln{p_{\theta}(\mathbf{x}|\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}})}\right]}_{\text{Reconstruction Loss}}-\lambda\underbrace{\mathbb{E}_{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}\sim q_{\alpha},q_{\eta},h_{\delta}}\left[\log q(\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}|\mathbf{x})-\log p(\mathbf{z})\right]}_{\text{KL Divergence}},$ (12)
where the reconstruction term ensures that the model accurately reconstructs the input data from the latent variables $\mathbf{z}_{c}$ and $\mathbf{z}_{m}$, while the KL divergence term regularizes the latent space to ensure the learned features align with the underlying data distribution. The hyperparameter $\lambda$ controls the trade-off between reconstruction accuracy and regularization strength.

### v2

@@ Abstract
Most existing studies on few-shot learning focus on unimodal settings, where models are trained to generalize to unseen data using a limited amount of labeled examples from a single modality. However, real-world data are inherently multi-modal, and such unimodal approaches limit the practical applications of few-shot learning. To bridge this gap, this paper introduces the Cross-modal Few-Shot Learning (CFSL) task, which aims to recognize instances across multiple modalities while relying on scarce labeled data. This task presents unique challenges compared to classical few-shot learning arising from the distinct visual attributes and structural disparities inherent to each modality. To tackle these challenges, we propose a Generative Transfer Learning (GTL) framework by simulating how humans abstract and generalize concepts. Specifically, the GTL jointly estimates the latent shared concept across modalities and the in-modality disturbance through a generative structure. Establishing the relationship between latent concepts and visual content among abundant unimodal data enables GTL to effectively transfer knowledge from unimodal to novel multimodal data, as humans did. Comprehensive experiments demonstrate that the GTL achieves state-of-the-art performance across seven multi-modal datasets across RGB-Sketch, RGB-Infrared, and RGB-Depth.
@@ 1 Introduction
Collecting large amounts of labeled data in real-world applications is often prohibitively expensive, time-consuming, or simply impractical [56, 53]. Few-shot learning (FSL) has emerged as a viable solution, enabling models to generalize effectively using only a handful of labeled examples [54, 5, 40, 23]. However, existing few-shot methods face significant challenges in handling the increasing prevalence of multi-modal data, such as multi-spectral images or multimedia content, which differs significantly from the extensive RGB data typically used in research. These multi-modal data collected from various types of sources and modalities, such as different sensors or imaging protocols, are becoming increasingly essential in applications like surveillance [66, 18] and medical image analysis [20, 46]. This highlights the need for more advanced FSL frameworks capable of leveraging the complementary information inherent in multi-modal data [40, 19].
@@
Figure 1: Comparison of recognition tasks. (a) Classical recognition requires extensive labeled data within a single modality. (b) Few-shot recognition uses a few labeled samples in a single modality to classify unseen samples. (c) Our proposed CFSL involves few labeled multi-modal samples and aims to generalize to unseen multi-modal samples from the same classes, leveraging both seen and unseen data from different modalities.
Recent efforts have leveraged large pre-trained foundation models to extend their capabilities on novel multi-modal tasks, including tabular data [71], audio [36, 10], and video classification [49], moving beyond traditional unimodal tasks like image classification [6]. Despite these advancements, the transfer of visual knowledge across different visual modalities remains relatively underexplored. Visual data, such as images and videos, constitute the most commonly studied data types; however, other visual modalities, such as infrared, depth, and sketches exhibit shared structural and contextual similarities with RGB data but possess unique attributes that make data collection and model adaptation more challenging.
In this case, this paper introduces a new Cross-modal Few-Shot Learning (CFSL) task designed to classify instances with multiple visual modalities with a limited number of labeled samples per class. In CFSL, the multi-modal data are organized into a support set and a query set. The support set comprises a few labeled examples from various modalities, while the query set contains unlabeled instances from corresponding classes and modalities that require classification. As illustrated in Figure 1, unlike classical supervised learning and FSL confined to a single modality, CFSL engages data from multiple visual modalities. The primary challenge of CFSL arises from the inherent variability and domain gaps across different visual modalities, which complicates feature extraction and alignment.
@@
Building on this understanding, previous studies [59, 60, 47] have highlighted the crucial role of compact visual representations and underlying concepts in enabling humans to recognize target objects across different contexts, which can be leveraged to enhance CFSL. As illustrated in Figure 2, the fundamental concepts like “bull” can be readily learned with only a few examples, thanks to the humans’ ability to generalize from limited data, regardless of visual modality [28, 4]. This capability, referred to as “evocation”, is grounded in the vast reservoir of previously encountered visual experiences that support recognition and generalization. Inspired by this observation, we hypothesize that models can achieve similar evocative capabilities by learning latent concepts from abundant unimodal data.
@@
To this end, we propose a Generative Transfer Learning (GTL) framework to facilitate knowledge transfer between unimodal and multi-modal data. In the GTL framework, we posit that the latent concepts underlying target objects consist of two components: (1) the intrinsic concept that captures the core characteristic transferable across modalities, and (2) the in-modality disturbance that accounts for variations unique to each modality. GTL aims to estimate these latent components and encode the relationships between the intrinsic concepts and visual content. Unlike existing methods that typically focus on modal adaptation [40, 41] or invariant feature extraction [77, 13] which fail to jointly model both variant and invariant components within a unified generative framework, GTL explicitly model the invariant components to be effectively transferred across modalities and the variant components to be adapted to specific modality characteristics, much like human cognitive processes.
Figure 2: Illustration of the ability to generalize concepts like “Bull” across visual modalities (Pablo Picasso. The Bull, 1945.).
@@
Within the GTL framework, our methodology consists of a two-stage training process. In the first stage, the generative learning stage, the model learns latent concepts label-free, relying solely on pre-trained visual representations. This stage focuses on capturing the intrinsic concept and the variations in visual content across modalities, ensuring that the learned relationship between the latent concept and visual content is robust and transferable. In the second stage, the recognition stage, the backbone network is frozen, and a separate classifier is trained on top of the learned latent intrinsic concepts to perform label prediction.
Our contributions are summarized as follows:
@@
  • We introduce a new cross-modal few-shot learning task, which focuses on the connection and distinction between visual modalities, requiring models to perform recognition on multi-modal data with minimal labeled samples. This task better reflects real-world scenarios, where multi-modal data is scarce and diverse.
  • We propose the generative transfer learning framework which explicitly models the variant and invariant components within a unified generative framework, enabling efficient knowledge adaption across modalities.
  • We demonstrate the effectiveness of our method through extensive experiments on seven multi-modal datasets across RGB-Sketch, RGB-Infrared, and RGB-Depth.
@@ 2 Task Settings
(a) t-SNE of feature distribution
of different modalities
(b) The data generating model
for the proposed CFSL
Figure 3: The observation of the severe modality differences and the details of the proposed generative model. (a) Illustration of the modality difference by the t-SNE clustering of the pre-trained CLIP [50] features of different modalities. (b) The proposed generative process for the representation learning stage, the green symbols are assumed to be parameters that enable the models to adapt from base to novel data.
@@
$D_{\text{support}}=\{(x^{i}_{m},y^{i})|m\in\{m^{1},m^{2},\dots,m^{d}\},y^{i}\in Y_{\text{novel}}\}$, a support set that contains a limited number of labeled samples per class from multiple modalities. Typically, this consists of $K$ labeled samples for each of $N$ classes.
$D_{\text{query}}=\{x^{j}_{m}|m\in\{m^{1},m^{2},\dots,m^{d}\}\}$, a query set contains massive of unlabeled samples from the same classes as $D_{\text{support}}$, but potentially from the different modalities.
The fundamental challenge in this task is the limited labeled data in $D_{\text{support}}$, combined with the multi-modal data in $D_{\text{novel}}$. The scarcity of labeled data per class, along with the need to generalize across different modalities, makes this task significantly more challenging.
As shown in Table 1, we provide a brief comparison between CFSL and existing tasks. For a more detailed discussion on related works and the distinctions between CFSL and existing tasks, please refer to Appendix A.
Data Phrase Properties CFSL(ours) FSL CD-FSL CM-FSL
Base - RGB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Other-Modal $\times$ $\times$ $\times$ $\times$
Novel tune RGB $\checkmark$ $\checkmark$ $\checkmark$/$\times$ $\times$
Other-Modal $\checkmark$ $\times$ $\times$$\checkmark$ $\checkmark$
test RGB $\checkmark$ $\checkmark$ $\checkmark$/$\times$ $\checkmark$
Other-Modal $\checkmark$ $\times$ $\times$$\checkmark$ $\times$
Table 1: Task comparisons. Unlike FSL [5, 40], CD-FSL (cross-domain) [14, 62, 68], and existing CM-FSL(cross-modality) [3], CFSL uniquely integrates multi-modalities in the novel phase, requiring consistent handling of RGB and other modalities.
@@
Figure 4: The proposed GTL framework. During the training on base data, all modules are trained (as in the blue dashed box), but when adapting to novel data, the generator is frozen, and all other parts are tunable (as in the red dashed box). The classifier for recognition is separately initialized on the base and novel training since there is no overlap on class between them.
@@ 3.1 Problem Formulation
$p(\mathbf{x},\mathbf{z},\mathbf{u},y)=p_{\delta}(\mathbf{z}_{m}|$ $\mathbf{u}^{1},\mathbf{u}^{2},\dots,\mathbf{u}^{d})p_{\gamma}(\mathbf{z}_{c}|y)$
$p_{\theta}(\mathbf{x}|\mathbf{z}_{m},\mathbf{z}_{c})p(\mathbf{u})p(y).$ (1)
@@
$\mathbf{z}_{m}\sim p_{\delta}(\mathbf{z}_{m}|\mathbf{u}^{1},\mathbf{u}^{2},\dots,\mathbf{u}^{d}).$ (4)
@@
$\mathbf{z}_{c}\sim p_{\gamma}(\mathbf{z}_{c}|y).$ (5)
@@
$\mathbf{x}\sim p_{\theta}(\mathbf{x}|\mathbf{z}_{m},\mathbf{z}_{c}).$ (6)
@@ 3.2 Network Design
In this section, we introduce the key components of our network, each playing a distinct role in modeling the cross-modal few-shot classification task. The specific details of the network architecture are presented in Appendix C. The operational framework of GTL is depicted in Figure 4, detailing the training and adaptation process for CFSL scenarios.
@@
$q_{\alpha}(\mathbf{{z}}_{c},\mathbf{{z}}_{m}|\mathbf{x})\sim\mathcal{N}(\mu,\sigma^{2}),$ (8)
@@
The in-modality disturbance latent variable $\mathbf{z_{m}}$ is assumed to be transferable when adapting from the base to novel data. Since there is a lack of direct supervision regarding which representation corresponds to modality-specific information, we employ a flexible estimator to approximate the unobservable prior $p(\mathbf{z}_{m})$. Inspired by the latent domain learning method [7], we use a set of learnable gating functions $g(\mathbf{x})$ that assign each observation $\mathbf{x}$ to multiple latent domains, capturing different perspectives.
@@
$\mathbf{\hat{z}}_{m}^{{}^{\prime}}=h_{\delta,\mathbf{\hat{u}}\sim q_{\eta}}(\mathbf{\hat{u}},\mathbf{\hat{z}}_{m})\quad\text{and}\quad q_{\eta}(\mathbf{\hat{u}}|\mathbf{x})=\sum_{d=1}^{D}g_{d}(\mathbf{x})V_{d}(\mathbf{x}),$
@@
The reconstruction module is responsible for generating an estimate of the observation $\mathbf{\hat{x}}$ based on the estimated latent variables $\mathbf{\hat{z}_{c}}$ (class-relevant information) and $\mathbf{\hat{z}_{m}^{{}^{\prime}}}$ (modality-relevant information). We adopt a generator with a similar but reversed structure as the posterior estimator. While the posterior estimator encodes the latent variables, the generator decodes them back to the observed data.
@@
$\hat{\mathbf{x}}=\text{Dropout}\left(\text{ReLU}\left(\text{MLP}(\mathbf{\hat{z}{{}^{\prime}}})\right)\right),\text{where}\quad\mathbf{\hat{z}}{{}^{\prime}}=\{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}\}.$
@@
The classifier $p_{\phi}(\mathbf{y}|\mathbf{z}_{c})$, parameterized by $\phi$, can be implemented using a simple linear classifier. This involves a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times c}$, where $d$ is the dimension of the latent variable $\mathbf{\hat{z}}_{c}$, and $c$ denotes the number of output classes. The classifier is trained by a standard cross-entropy loss function. The linear classifier computes the logits by linearly combining the latent variable $\mathbf{\hat{z}}_{c}$ with the weight matrix $\mathbf{W}$, followed by a softmax function to output class probabilities. The logits are computed as:
@@ 3.3 Learning Objectives
$\mathcal{L}_{\text{ELBO}}=\underbrace{\mathbb{E}_{q_{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}|\mathbf{x}}}\left[\ln{p_{\theta}(\mathbf{x}|\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}})}\right]}_{\text{Reconstruction Loss}}$
$-\lambda\underbrace{\mathbb{E}_{\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}\sim q_{\alpha},q_{\eta},h_{\delta}}\left[\log q(\mathbf{\hat{z}}_{c},\mathbf{\hat{z}}_{m}^{{}^{\prime}}|\mathbf{x})-\log p(\mathbf{z})\right]}_{\text{KL Divergence}},$ (12)
where the reconstruction term ensures that the model accurately reconstructs the input data from the latent variables $\mathbf{z}_{c}$ and $\mathbf{z}_{m}$, while the KL term regularizes the latent space to ensure the learned features align with the underlying data distribution. The hyperparameter $\lambda$ controls the trade-off between reconstruction accuracy and regularization strength.